0.0075mol of CaCO3 is added to 0.015mol of a solution of hcl, what is the volume of gas evolved at S.T.P
1
Expert's answer
2017-02-22T09:19:06-0500
CaCO3 + 2HCl = CaCl2 + H2O + CO2
V = Vm*n
V(CO2) = 0.0075. mol * 22.4 L/mol = 0.168 L
Answer: 0.168 L